# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v1

@@ 5.3. FINANCIAL DISCIPLINE AND UPI USAGE
Our research shows the pivotal role of financial discipline in UPI usage. User interviews revealed a distinct contrast between those with strong financial discipline, harnessing UPI’s convenience without compromising financial stability, and those who occasionally succumbed to overspending. The latter group often cited a lack of financial literacy and self-control contributing to their spending behavior. The abstract nature of digital transactions in UPI can sometimes lead to impulsive purchases. Therefore, it is imperative to promote financial education and in helping the user understand their expenses and responsible spending practices alongside adopting digital payment systems like UPI. This two-pronged approach can empower users to maximize the advantages of UPI while safeguarding their financial well-being.
5.4. FUTURE IMPLICATIONS AND POLICY CONSIDERATION
The research outcomes hold significant implications for various stakeholders in India. Policymakers can use this information to design regulations that encourage responsible UPI usage and protect users from potential scams. Businesses and financial institutions should adapt their services to accommodate the diverse needs of users from various age groups and professions, enhancing their overall user experience. Additionally, educational initiatives can focus on improving financial literacy and promoting responsible spending habits among UPI users. This research emphasizes the transformative power of UPI in reshaping spending behavior, making it crucial for all involved parties to adapt and harness its full potential for the benefit of Indian consumers and the broader economy. Many other countries, like France, Singapore, etc., have started embracing UPI to promote digital payments (glo, 2023). With the recent developments by NPCI like UPI integration with Rupay Credit Card (rup, 2012), tap and pay feature with Near Field Communication (NFC) (nfc, 2004), UPI Lite (upi, 2022) and UPI LiteX (upi, 2023) for offline payments, the infrastructure of UPI is being aimed to scale to 100 billion monthly transactions, making it a greater force of revolution in the Indian financial landscape.

### v2

@@ 5.3. FINANCIAL DISCIPLINE AND UPI USAGE
Our research highlights the importance of financial discipline in UPI usage. Users with strong financial discipline leverage UPI for convenience and stability, while those prone to overspending often cite a lack of financial literacy. The abstract nature of UPI transactions can lead to impulsive purchases. Promoting financial education is essential to address this, empowering users to understand expenses and practice responsible spending while embracing UPI. This dual approach empowers users to benefit from UPI while safeguarding their financial well-being.
5.4. FUTURE IMPLICATIONS AND POLICY CONSIDERATIONS
The research outcomes have significant implications for stakeholders in India. Policymakers can design regulations promoting responsible UPI usage and protecting users from scams. Businesses and financial institutions should tailor services to diverse user needs, enhancing overall user experience. Educational initiatives can focus on improving financial literacy and responsible spending habits among UPI users. This underscores the transformative power of UPI, urging all parties to adapt for the benefit of Indian consumers and the broader economy. Many other countries, like France, Singapore, etc., have started embracing UPI to promote digital payments (glo, 2023). With the recent developments by NPCI like UPI integration with Rupay Credit Card (rup, 2012), tap and pay feature with Near Field Communication (NFC) (nfc, 2004), UPI Lite (upi, 2022) and UPI LiteX (upi, 2023) for offline payments, the infrastructure of UPI is being aimed to scale to 100 billion monthly transactions, making it a greater force of revolution in the Indian financial landscape.
@@ 6. Conclusion
Appendix A SURVEY QUESTIONS
1. What is your age?
2. What is your profession?
3. What is your gender? (Options: Male, Female, Non-binary, Prefer not to say)
4. Have you used UPI as the payment method?
5. If yes, which year did you start using it?
6. How often do you use UPI for making payments? (Daily, Weekly, Monthly, Occasionally, Rarely)
7. Have you noticed any changes in your spending habits since you started using UPI?
8. On a scale of 1 to 5, rate the convenience you find with UPI for making payments.
9. How satisfied are you with your overall experience using UPI for payments? (1 = Not Satisfied at all, 5 = Extremely Satisfied)
10. On average, how much do you spend daily using UPI?
11. Have you found any change in your spending ever since using UPI?
12. Do you feel that using UPI affects your ability to stick to your budget?
13. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage?
14. How often do you check your transaction history or keep track of your UPI spending? (options: rarely, occasionally, frequently, very frequently, regularly)
15. How do you manage/track your expenses? (Which app do you use, if you do, e.g., split-wise, etc)
16. Additional Comments: Do you want to share anything about your experience with UPI and its impact on your spending behavior?
17. Will you be available for online interviews if called? (Yes/No)
18. If yes, provide your email address :
19. If yes, provide your phone number :
Appendix B INTERVIEW QUESTIONS
Demographic Information:
a. What is your age?
b. What is your gender? (Options: Male, Female, Non-binary, Prefer not to say)
c. What is your occupation?
Usage of UPI:
a. Have you used UPI as a payment method? (Yes/No)
b. If yes, when did you start using it?
c. How often do you use UPI to make payments? (Daily, Weekly, Monthly, Occasionally, Rarely)
d. Can you briefly describe your experience with using UPI to make transactions?
e. Have you noticed any changes in your spending habits since you started using UPI? (options: ”Increased spending,” ”Decreased spending,” or ”No change”)
f. If yes, could you elaborate on the changes you have observed in your spending habits after adopting UPI?
Perception of UPI Convenience:
a. On a scale of 1 to 5, rate the convenience you find with UPI for making payments. (1 = Not Convenient at all, 5 = Extremely Convenient)
b. In what ways do you perceive UPI to be convenient/inconvenient for you?
c. On a scale of 1 to 5, how would you rate the convenience of using UPI compared to traditional payment methods like cash? (1 = Not Convenient at all, 5 = Extremely Convenient)
d. How satisfied are you with your overall experience of using UPI for payments? (1 = Not Satisfied at all, 5 = Extremely Satisfied)
Income Information:
a. What is your approximate annual income range?
Spending Behavior:
a. On average, what is the transaction amount when you use UPI?
b. On average, how much do you spend daily using UPI?
c. On average, how much do you spend monthly using UPI?
c1. Where do you spend your money?
c2. Does the spending amount include only UPI transactions or all spending in general?
d. Have you found any change in your spending ever since using UPI?
e. If yes, could you specify the reasons for the increased spending after adopting UPI?
f. Do you feel that using UPI affects your ability to stick to your budget? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
g. How often do you check your transaction history or keep track of your UPI spending? (Options: Rarely, Occasionally, Frequently, Very Requently, Regularly)
h. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage?
i. In your opinion, does using UPI encourage you to spend more money overall? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Factors Influencing Spending Behavior:
a. Besides UPI usage, do other factors have influenced your spending behavior?
b. If yes, please specify those factors.
c. Are you aware of any psychological factors influencing your spending behavior when using UPI? If yes, please elaborate.
Attitudes Towards Financial Management:
a. How would you describe your level of financial discipline? (1 to 5)
b. Do you feel that using UPI has made managing your finances easier or more difficult? (“Easier,” “More Difficult,” “No Change”)
c. How do you rate the statement ”Do you feel that UPI provides you with a clear understanding of your spending patterns”? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
d. How do you rate the statement ”UPI has simplified your financial management”? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Additional Comments:
a. Is there anything else you would like to share about your experience with UPI and its impact on your spending behavior?